4. In mice, the Ayallele of agouti is dominant for yellow coat color but is homozygous lethal. A yellow mouse, Ay/A, that is also heterozygous for the albino allele, C/c, is crossed to an albino mouse that is also heterozygous for the Ayallele, Ay/A. c/c. What phenotypic ratios do you expect in the progeny? (True or False)

a. if these genes are unlinked: 3/8 yellow, 1/8 agouti, 1/2 albino

b. if these genes are unlinked: 1/3 yellow, 1/6 agouti, 1/2 albino
c. if these genes are very closely linked: 9/16 yellow, 1/16 agouti, 6/16 albino

d. if these genes are very closely linked: 2/3 yellow, 1/3 agouti
e. whether or not the genes are linked, 1/2 of the total possible progeny die because of the Ay/Ay lethality

5. Two wild cabbage-related plant strains form an allotetraploid plant. The two diploid parent strains have the genotypes A/A and a/a at the same gene on their respective chromosome 1. What ratios of progeny genotypes would be expected for an allotetraploid plant, under the following conditions: (True of False)

a. crossed back to the a/a parental strain= 1A/A/A/A: 1A/A/a/a

b. crossed back to the a/a parental strain= 3 A/A/A: 1A/A/a

c. self-fertilized = 1 A/A/A/A: 2 A/A/a/a: 1 a/a/a/a

d. self-fertilized = all A/A/a/a

e. self-fertilized = 3 A/A/A/A: 1 A/A/a/a

6. The Xg blood group is the only blood group system located on the X chromosome. There are two alleles: one that makes antigen ("a"), which is dominant over the other allele, which makes no antigen ("o"). A type "a" man and type "o" woman have a son with Klinefelter's syndrome (XXY), who is type "a". What can you conclude about this child's disorder? (True or False)

a. the nondisjunction could have happened during meiosis I in the mother's germline

b. the nondisjunction could have happened during meiosis II in the mother's germline

c. the nondisjunction could have happened during meiosis I in the father's germline

d. the nondisjunction could have happened during meiosis II in the father's germline
e. the child’s X chromosomes are unlikely to be homozygous at the Xg locus
Answer #1

AYACc * AYAcc

AYC AYc AC Ac
AYc AYAYCc (lethal) AYAYcc (lethal) AYACc (yellow) AYAcc (albino)
Ac AYACc (yellow) AYAcc (albino) AACc (agouti) AAcc (albino)

if the genes are unlinked 1/3 (2/6=1/3) yellow,1/6 agouti, 1/2 (3/6=1/2) albino will be produced. ( here total number of progeny=6)

If the genes are closely linked then these 2 genes moves tother so

AYACc ( assuming alleles are in cis, that is both dominant alleles AY and C are on one chromosome and the recessive alleles A and c are on the other homologous chromosome ) can produce only AYC and Ac gametes

and the other parent AYAcc can produce only AYc and Ac gametes

so the cross is

AYACc * AYAcc

AYC Ac
AYc AYAYCc (lethal) AYAcc (albino)
Ac AYACc (yellow) AAcc (albino)

if the genes are tightly linked

1/3 yellow and 2/3 ( 2 albino out of 3 ) are produced.

so

a. if these genes are unlinked: 3/8 yellow, 1/8 agouti, 1/2 albino- False

b. if these genes are unlinked: 1/3 yellow, 1/6 agouti, 1/2 albino- true
c. if these genes are very closely linked: 9/16 yellow, 1/16 agouti, 6/16 albino- False

d. if these genes are very closely linked: 2/3 yellow, 1/3 agouti- false
e. whether or not the genes are linked, 1/2 of the total possible progeny die because of the Ay/Ay lethality- false